OpenStudy (anonymous):

given that f(x)=x^2+3x-10 and g(x)=4x+3 find (g o f)(1)

OpenStudy (anonymous):

g o f(x) = 4(x^2 + 3x - 10 )+3 g o f(1) = 4(1^2 +3(1) -10) +3 = -21